### v1

# Sensory Glove-Based Surgical Robot User Interface

Abstract
Robotic surgery has reached a high level of maturity and has become an integral part of standard surgical care. However, existing surgeon consoles are bulky and take up valuable space in the operating room, present challenges for surgical team coordination, and their proprietary nature makes it difficult to take advantage of recent technological advances, especially in virtual and augmented reality. One potential area for further improvement is the integration of modern sensory gloves into robotic platforms, allowing surgeons to control robotic arms directly with their hand movements intuitively. We propose one such system that combines an HTC Vive tracker, a Manus Meta Prime 3 XR sensory glove, and God Vision wireless smart glasses. The system controls one arm of a da Vinci surgical robot. In addition to moving the arm, the surgeon can use fingers to control the end-effector of the surgical instrument. Hand gestures are used to implement clutching and similar functions. In particular, we introduce clutching of the instrument orientation, a functionality not available in the da Vinci system. The vibrotactile elements of the glove are used to provide feedback to the user when gesture commands are invoked. A preliminary evaluation of the system shows that it has excellent tracking accuracy and allows surgeons to efficiently perform common surgical training tasks with minimal practice with the new interface; this suggests that the interface is highly intuitive. The proposed system is inexpensive, allows rapid prototyping, and opens opportunities for further innovations in the design of surgical robot interfaces.
I INTRODUCTION
Robotic surgery has reached a high level of maturity and has become an integral part of standard surgical care. For the surgeon, it offers increased precision and dexterity through efficient interaction interfaces. However, despite its high acceptance and success, it faces some challenges in modern operating rooms. The surgeon consoles are bulky and take up valuable space [1]. A redesigned workspace promises to improve turnaround time, increase throughput, and reduce cost [2]. Studies further emphasize the potential of ergonomic workspace design and nonoperative process redesign to advance the efficiency of robotic surgical devices [3, 4]. Another challenge with current consoles is that the surgeon is separated from the rest of the team, which can lead to challenges in communication [5] and can impact the efficiency of the operational workflow in tasks such as robotic docking [6]. Finally, the proprietary nature of existing consoles makes it difficult to innovate in this space despite rapid advances in fields such as augmented and virtual reality that could benefit the surgeon.
Several studies have attempted to address some of these shortcomings. [7] proposes a novel method for controlling the endoscope through a head-mounted master interface (HMI) integrated with a surgical robot system. [8] controls the surgical robot arm and the Augmented Reality (AR) interface through various hand gestures. In [5], the authors provide a nice review of existing approaches and opportunities.
One potential area for further improvement and technological innovation in robotic surgery is the integration of modern sensory gloves into robotic platforms, which allows surgeons to intuitively control robotic arms with their hand movements without the need for additional manipulators that are customary for existing consoles. Furthermore, most end-effectors of surgical robot instruments, such as the Fenestrated Bipolar Forceps (FBF) and the Large Needle Driver (LND), are limited to the use of opposing blades due to the design of the console manipulators (as the last joint is controlled only by pinching two fingers); glove-based control could offer the surgeon additional dexterity while using these instruments.
Glove-based control is intuitive and is popular in robotics applications outside the surgical field. For instance, a low-cost wireless system with a glove equipped with bending sensors is described in [9]. In [10], the authors demonstrate glove-based control in rehabilitation, using gloves equipped with Inertial Measurement Units (IMUs) for patient exercises. Glove-based control for telemanipulation tasks, integrating force sensors for enhanced feedback, is explored in [11]. These studies explore various sensor modalities and develop communication techniques and control algorithms to achieve sophisticated and effective glove-based control systems for robotic arms.
In our work, we develop a system in which an HTC Vive Tracker [12] is attached to a Manus Meta Prime 3 Haptic XR sensory glove [13], equipped with multiple IMUs and bending sensors for finger tracking, to control one arm of the da Vinci surgical robot paired with the da Vinci Research Kit (dVRK) [14]; the visual feedback to the surgeon is provided through God Vision wireless smart glasses [15]. In addition, this setup manipulates the robotic arm’s end effector (jaw of the instrument) through finger movements. Moreover, we implemented hand gesture recognition to substitute pedal inputs of the current da Vinci console (e.g. clutch). This also allows us to clutch the instrument orientation, a functionality not provided by the da Vinci system. The glove is equipped with vibrotactile elements to provide feedback to the user about whether the gesture is activated.
We evaluate the proposed system in terms of tracking accuracy and the ability of the user to perform surgical training tasks similar to those used in training with the da Vinci devices. The evaluation shows that the system is responsive and that surgeons can effectively complete training exercises with minimal practice with the new interface, suggesting that the interface is highly intuitive. The proposed system is inexpensive as it uses off-the-shelf components and it dramatically reduces space requirements. It demonstrates how modalities such as vibrotactile feedback can efficiently provide additional information to the surgeon without cluttering the visual field. Finally, the proposed architecture allows for rapid prototyping and opens opportunities for further innovations in the design of surgical robot interfaces.
II METHODOLOGY
Figure 1: The setup showing the configuration of the da Vinci arm (PSM) and the Manus glove with the HTC Vive Tracker attached. The user wears the God Vision smart glasses for the stereo endoscopic view. Shown are the $x$ (red), $y$ (green), and $z$ (blue) axes of the relevant coordinate frames.
II-A System Architecture
The system integrates several components. The surgical robot platform consists of the first-generation da Vinci robot paired with the dVRK. One of the Patient Side Manipulators (PSMs) is equipped with the Fenestrated Bipolar Forceps (FBF) instrument and is subject to glove control. The arm is configured as in Fig. 1. The HTC Vive Tracker attached to the Manus glove provides the pose of the hand, while the fingers are tracked by the Manus Meta Prime 3 XR. In addition, classifier models were trained for hand gesture recognition; hand gestures were used to supersede da Vinci’s pedal inputs. The visual feedback to the user is provided through God Vision smart glasses. The system runs the Robot Operating System (ROS) [16] that facilitates seamless communication and coordination between the Manus glove, the HTC Vive Tracker, and dVRK, ensuring real-time data streaming and responsive interaction between the user and the control interface. The architecture allows for easy prototyping.
II-B da Vinci Arm Kinematics
For the full da Vinci system, the kinematic chain from the Patient Side Manipulator (PSM) tip to the Endoscope Camera Manipulator (ECM) tip involves Setup Joints (SUJs). The poses provided by the dVRK exhibit errors within a range of ±$\pm 5cm$ for translation and fall between $5\sim 10$ degrees for orientation These inaccuracies arise due to challenges in calibrating the SUJs precisely, as reported in [14]. Consequently, we implemented a custom calibration method for the dVRK using fiducial markers [17] that eliminates the influence of SUJs and allows us to accurately determine the transformation between the PSM base (frame $R$) and the instrument tip (frame $T$) (Fig. 1).
II-C Glove Kinematics
To monitor the glove’s position and orientation within the environment, we attached an HTC Vive Tracker to it. The glove’s position and orientation data are updated at 120 Hz using the data from the IMUs and HTC Vive Base Stations [18, 19].
The coordinate frames for the HTC Vive Tracker and Base Station are depicted in Fig. 1. At initialization, the tracker frame $H$ and the tracker base frame $S$ coincide and are aligned with the PSM instrument tip frame $T$. A displacement of $H$ from $S$ results in the scaled motion of PSM (see below) to create an intuitive control for the user.
II-D Control Architecture
To ensure safety for both the glove and the instrument tip, we define their operational workspaces to be cubes, with a larger side $L_{h}$ for the glove and a side $L_{t}$ for the instrument tip, centered at their initial positions. In both cases, movements beyond the cube volume are projected to the cube surface.
The movements of the glove tend to be significantly larger than the desired movements of the PSM tip ($L_{h}>L_{t}$). The motion of the glove is thus scaled to match the range needed for PSM operations:
$\Delta p_{t}=\eta\,\Delta p_{h}$ (1)
where $\Delta p_{h}$ represents the translational movement of the glove from its initial position, and $\Delta p_{t}$ denotes the expected translational movement of the PSM from its initial position, and $\eta={L_{t}}/{L_{h}}$ is the scaling factor. Similarly to the existing surgeon console, the scaling factor $\eta$ can be adjusted.
II-E Finger Tracking
The tracking of the relative transformation between the fingers and the wrist is accomplished using the five sets of flexible sensors with 2 Degrees of Freedom (DoF) and six sets of 9 DoF IMUs of the Manus Meta Prime 3 Haptic XR glove, operating at a rate of 90Hz. These sensors boast a ±$\pm 2.5^{\circ}$ accuracy [13]. The glove software provides a hand skeleton mesh, inspired by the Mediapipe framework [20]. This mesh is dynamically adjusted based on the readings from the glove sensors, allowing for a real-time representation of the hand pose.
In our architecture, the Euclidean distance between the tips of the index and thumb fingers is used to control the jaw of the end effector. To map the distance of the fingers to the distance between the tips of the jaw we use an expression analogous to Eq. (1). This enables precise and intuitive control of the robotic system.
II-F Hand Gesture Recognition
Hand gestures serve as a seamless and intuitive means of interaction and open up the possibility of introducing a large variety of inputs that can surpass what the current consoles can provide. For example, in [8] it was shown that it is possible to implement 8 different commands.
In our work, we train 4 different multiclassification models for hand gesture recognition using our custom dataset. Our implemented architecture can currently distinguish between five distinct gestures: None, Pinky (thumb and pinky in contact), Ring (ring finger and thumb in contact), Fist, and Thumb up, as illustrated in Table I. Our framework allows a new gesture to be added in just a few minutes. Currently, only three of the five gestures serve as commands (Fist, Ring, and Pinky). Fist activates the clutch mechanism, Ring, if held for 2 seconds, initiates and suspends robot tracking, and Pinky triggers monopolar current delivery [21]. See Section III for further details.
II-G Clutch
The clutch functionality in the da Vinci surgical robot serves as a safety feature and a means of control for the surgeon during robotic-assisted surgeries. This mechanism essentially allows the surgeon to recenter the workspace of the surgeon console by temporarily disengaging the movement of the robotic arms while still maintaining control of the Master Tool Manipulators.
When the surgeon operates the da Vinci system’s master console, which consists of hand and finger controls, they can engage or disengage the clutch as needed. This allows them to stop the movement of the robotic arms, enabling them to make adjustments or reposition the controllers as required during the procedure.
In the dVRK system, the clutch is engaged with a foot pedal or a button accessible by the index finger. However, the current rotation of the tool is maintained and cannot be changed (this is achieved by blocking the last two joints of the controller). In contrast, our system implements full clutching (both translation and rotation). The clutch is engaged by closing the hand (Fist gesture) as described next.
The clutching mechanism includes 3 phases:
Leading Edge: When the Fist gesture is detected, the clutch is engaged, and the vibrotactile feedback of the glove is activated. The current configuration of the PSM $g_{rt}^{0}$ is recorded and the robot is prevented from moving. Note that the jaw of the instrument is allowed to move.
Active Clutch: The vibrotactile feedback continues to be activated. The PSM is held immobile. The glove can be moved to an arbitrary pose, allowing the user to move to a comfortable configuration.
Trailing Edge: When the gesture becomes other than Fist, the clutch is disengaged and the vibrotactile feedback is deactivated. The tracker frame $H$ and the tracker base frame $S$ are aligned, and their axes are aligned with the axes of the PSM tip frame $T$ whose configuration is described by $g_{rt}^{0}$.
III Hand Gesture Recognition
Hand Gesture Representation Hand Gesture Representation
Thumbs up [Uncaptioned image] Pinky [Uncaptioned image]
Ring [Uncaptioned image] Fist [Uncaptioned image]
TABLE I: Gestures trained for distinct functionalities. In this paper, Fist invokes the clutch mechanism and Ring toggles the robot control on or off. None gesture corresponds to the case when the subject is not using any of the four gestures above.
To train the classifiers for hand gesture recognition, we created a custom dataset. We used 5 sets of approximately $3-5$ minute recordings from the glove worn by one of the authors while performing one of the 4 gestures reported in Table I, or none of them. The recorded data (Table II) consists of the point cloud of each of the 21 landmark relative poses (3 position entries and 4 quaternion entries) regarding the wrist IMU (147 features). Table III shows examples of point clouds for different gestures. Each recording has been made while changing the position and orientation of the hand as much as possible to ensure the widest variety of data. Each recording corresponds to one gesture (or no gesture) so when merging all the data it was easy to attribute a class to each sample. As the number of classes is small, a simple one-hot encoding has been used.
Hand Gesture None Pinky Ring Fist Thumbs up
Train set $20743$ $12833$ $25011$ $16738$ $19447$
Test set $4167$ $2493$ $5075$ $3445$ $3775$
TABLE II: The size of our training and testing sets for hand gesture recognition.
Input Hand Pose Predicted Gesture Input Hand Pose Predicted Gesture
[Uncaptioned image] Thumbs-up [Uncaptioned image] Pinky
[Uncaptioned image] Ring [Uncaptioned image] Fist
TABLE III: Examples of point clouds used for gesture prediction.
III-A Classifiers
We initially used Random Forests (RF) [22] as a model for gesture recognition. Subsequently, we explored the Multilayer Perceptron (MLP) classifier [23]. The MLP model was configured with two hidden layers containing 40 and 25 neurons, respectively, and a maximum of 100 iterations. The use of such a small model with only two hidden layers helps prevent overfitting.
In addition to these established models, we investigated a novel classifier, the Light Gradient Boosting Machine (L-GBM) [24]. L-GBM employs a sequential approach in which decision trees are added successively, with each tree correcting errors from the ensemble of preceding trees. Unlike traditional gradient-boosting methods, L-GBM directly minimizes the loss function. Tailored for efficient training on large datasets, it is particularly suitable for applications with resource constraints, and it can quickly predict data.
We also trained a Histogram-based Gradient Boosting model that combines the traditional gradient boosting [25] and the histogram-based algorithm [24]. Unlike traditional gradient boosting methods that operate on individual data points, histogram-based gradient boosting works with binned feature values. By constructing histograms of input features, this approach accelerates the computation of information gained during tree construction. The binning process reduces the number of unique feature values, resulting in faster training times and decreased memory usage.
III-B Results
By training the models above, it is possible to recognize 5 different gestures (tested on a new, unseen dataset). The models were implemented in ROS and Table IV shows their performance. The confusion matrices for each model are shown in Fig. 2.
Model Accuracy F1 Score Recall
MLP 0.9852 0.9852 0.9852
L-GBM 0.9508 0.9505 0.9508
RF 0.9417 0.9413 0.9417
HGBT 0.9628 0.9627 0.9628
TABLE IV: Evaluation metrics for the trained models.
Although all models exhibited high accuracy, F1 scores, and recall, we selected the MLP model for implementation due to its high prediction speed (Table V) and high accuracy. The average frequency at which the MLP model can operate (determined from the ROS logs) is around 96Hz. From Fig. 2, it is noticeable that the Pinky and Ring gestures are associated with more errors. This observation can be attributed to the similarity between these two gestures.
(a) MLP
(b) L-GBM
(c) Random Forest
(d) HGBT
Figure 2: Confusion matrices for gesture recognition models.
Model Mean Time (ms) Std. Deviation (ms)
MLP 0.75 0.14
L-GBM 2.09 3.06
RF 0.92 0.16
HGBT 7.81 4.72
TABLE V: Prediction time for each model (1000 trials).
In our application, gesture recognition needs to be performed continuously and in real-time. To ensure the stability of recognition – for example, even a single mistake in the clutch could lead to problematic input sent to the robot – we use a sliding window of length 7. The predictions of the MLP model over the last 7 time steps are stored in a 7x5 matrix, with each row corresponding to a time sample and each column representing a gesture prediction (0 or 1). The matrix is initialized with zeros. During each iteration, a new row is added with the most probable gesture set to one; the oldest prediction is deleted from the matrix. Finally, in each iteration, the system selects the gesture with the highest column sum. This induces a negligible delay but greatly increases the stability of the prediction.
IV Preliminary Evaluation
To evaluate the system, the members of the team with surgical training were asked to use the system to complete one of the standard da Vinci training exercises [26]. The original exercise was modified to suit our system’s single-arm setup and to simplify the task, thereby reducing any uncertainties in the preliminary test. The modified task required participants to control the PSM with their left hand to lift a peg from one pin on the pegboard and transfer it to another pin. One trial consists of repeating this task twice with two different pegs on the pegboard.
IV-A Experiment Setup
The experimental setup is shown in Fig. 1. The user’s hand starts in a position similar to that shown in the figure, away from the robot. The endoscope was positioned to allow a view of the pegs on the board and the tip of the instrument. God Vision glasses were connected to the vision tower, and the user wore the glasses to access the stereo endoscopic view. The evaluation was carried out by 5 surgeons and one fellow, all members of the research team. All have prior experience with the da Vinci surgical console to perform surgical tasks. Participants underwent a brief training session, lasting approximately 2 to 10 minutes, primarily to familiarize themselves with the operation of the clutch mechanism. Each participant performed the exercise 5 times. To implement our system, we needed to update the dVRK. This caused compatibility issues with the existing dVRK console so we were not able to conduct the exercise through the standard da Vinci console.
IV-B Control Results
(a) Translation
(b) Orientation
(c) Jaw Movements
Figure 3: A sample recording of motion between the glove and the PSM. The segments highlighted in purple indicate periods when the clutch mechanism was engaged. (a) displays the translation of the glove and PSM while (b) shows the orientation of the glove and PSM respectively. (c) compares the distance of the thumb and index finger on the glove with the jaw angle of the PSM.
Figure 4: Translational and orientational distance between the glove and the PSM: the left figure depicts the Euclidean distance, while the right one details the orientation difference, quantified with rotation matrices (in the hand workspace).
Figure 5: Both plots detail the x-axis translation of the glove and the PSM tip in Fig. 2(a). The initial plot illustrates a delay in the robot’s execution of the glove’s movements. To address this, the delay between the signals was measured based on the peaks, and the aligned signals were shown on the bottom.
A sample trajectory of the da Vinci arm controlled with our proposed glove, performed by one of the surgeons, is shown in Fig. 3. A comparison of the kinematics of the glove and the instrument tip is shown in Figs. 2(a) and 2(b), respectively. The motion of the PSM exhibits a stair-case movement due to the implementation of position control of the arm in dVRK (green box in Fig. 5), where the robot temporarily stops once it reaches the goal position. The shaded regions show the moments when the clutch (Fist gesture) was activated. Our clutch mechanism (Section II-G) prevents the PSM from moving while activated and resumes the movement smoothly when released. The motion of the index and thumb of the glove and the instrument jaw is depicted in Fig. 2(c). Currently, the fully closed configuration of the FBF jaw is represented by a negative value in the dVRK’s system (as the $0^{\circ}$ angle corresponds to a parallel configuration of the jaw where it is not completely closed); the slight difference for extreme values between the glove and the PSM is due to the scaling between the glove and the jaw.
To compare the input motion of the glove and the output motion of the instrument tip, we first reversed the scaling (the inverse of Eq. (1)) of the movement of the instrument tip (Fig. 5). The figure clearly shows a delay between the two trajectories, mainly due to the computation time for the inverse kinematics. To accurately estimate the error between the input and output signals and eliminate the effect of the delay, we identified the delay by the peak positions within each signal. Subsequently, the PSM signal is shifted according to the calculated delay to align with the glove’s signal (Fig. 5). The difference in translation was quantified using the Euclidean distance, while the discrepancy in rotation was determined using the standard bi-invariant metric on the group of rotations [27]. The errors derived from Figs. 2(a) and 2(b) are shown in Fig. 4. The duration and error metrics for each participant’s performance are depicted in Table VI. The average delay in all trials and all participants was 0.223s.
User Average Duration (s)
Translational
Error (m)
Rotational
Error (rad)
Average Std. dev. Average Std. dev.
A 60.345 0.005 0.004 0.024 0.024
B 45.304 0.004 0.004 0.016 0.018
C 57.922 0.005 0.004 0.040 0.056
D 53.298 0.004 0.004 0.021 0.026
E 52.506 0.003 0.003 0.015 0.015
F 37.782 0.003 0.003 0.021 0.021
TABLE VI: the first column displays the average runtime for the five trials. The subsequent columns detail the overall mean and standard deviation of the errors, calculated based on the glove’s workspace, performed by each participant (complete version available on [28]).
IV-C Survey Results
In the follow-up survey, each participant was asked five questions: 1) How comfortable was the glove controller? 2) How responsive was the robot when controlled with the glove controller? 3) How intuitive was controlling the jaw with your fingers? 4) How intuitive was it to control the arm with the glove? 5) Was the vibrotactile feedback for the clutch effective? We use a 7-point Likert scale for the response, with 7 being the most positive response [29]. The average responses for the 6 participants are summarized in Table VII. The results of Questions 3 to 5 show that the new control architecture is intuitive. This is further evidenced by the small amount of practice required to complete the exercise. Questions 1 and 2 produced satisfactory results; all participants responded that the system was ’highly responsive.’ However, the results of Question 1 indicate that the overall ergonomics of the system can be improved, as most users had to lift their elbow to achieve a correct angle with their hand. This issue also highlights two important aspects: first, the user’s experience with the device may significantly impact their comfort, as they can readjust themselves by clutching to find a better position. Second, an ergonomic seat can greatly reduce the user’s effort.
Question # Q1 Q2 Q3 Q4 Q5
Avg. Score 5.50 6.00 6.50 6.50 6.33
TABLE VII: Average scores in the Likert scale from the users.
V Conclusion
This paper presents a novel user interface for the da Vinci surgical robot arm. The system integrates the Manus Meta Prime 3 Haptic XR glove, the HTC Vive Tracker, God Vision wireless smart glasses, and controls one arm of the dVRK surgical robot platform. In addition to moving the arm, the surgeon can use fingers to control the end-effector of the surgical instrument. A custom robot calibration was implemented to address the inherent inaccuracies of the dVRK system. A combination of scaling and gesture recognition was used to achieve precise and intuitive control of the robot’s movements while implementing an upgraded version of the clutch mechanism. In particular, we introduced clutching of the instrument orientation, a functionality not available in the da Vinci system. The vibrotactile elements of the glove are used to provide feedback to the user when gesture commands are invoked.
The proposed interface offers several potential advantages. Surgeons can directly control robotic arms by moving their hands, without the need for additional physical devices. This facilitates a more immersive surgeon-robot interaction. However, it should be noted that the hardware used in our setup is not the latest available on the market; we used the first version of HTC Vive Trackers and Manus Meta Prime 3 XR, both of which have lower precision compared to their upgraded versions (HTC Vive Trackers v2 and Quantum XR Metaglove).
The system was tested by members of the team with surgical training, including 5 surgeons and 1 fellow. They were asked to perform a surgical training task similar to those used in training with the da Vinci devices. The evaluation shows that the system is responsive and precise. Surgeons were able to effectively complete training exercises with minimal practice with the new interface, suggesting that the interface is highly intuitive. After the exercise, each participant filled out a questionnaire to provide feedback on their user experience. The responses show a high level of user satisfaction.
Existing surgical consoles are bulky and take up valuable space. They limit the ability to optimize the operating room setup for throughput. They also pose challenges to the communication of the surgical team, leading to potential workflow disruptions. In contrast, our proposed system is inexpensive as it uses off-the-shelf components, and it dramatically reduces space requirements. It demonstrates how modalities such as vibrotactile feedback can efficiently provide additional information to the surgeon without cluttering the visual field. Finally, the proposed architecture allows for rapid prototyping and opens opportunities for further innovations in the design of surgical robot interfaces.
Future work includes implementing velocity control to enhance user experience, upgrading hardware to improve tracking precision, enabling full control for the right hand, and expanding the repertoire of user gestures to increase the range of available actions.

### v2

@@
 Robotic surgery has reached a high level of maturity and has become an integral part of standard surgical care.
-However, existing surgeon consoles are bulky and take up valuable space in the operating room, present challenges for surgical team coordination, and their proprietary nature makes it difficult to take advantage of recent technological advances, especially in virtual and augmented reality.
-One potential area for further improvement is the integration of modern sensory gloves into robotic platforms, allowing surgeons to control robotic arms directly with their hand movements intuitively.
-We propose one such system that combines an HTC Vive tracker, a Manus Meta Prime 3 XR sensory glove, and God Vision wireless smart glasses.
+However, existing surgeon consoles are bulky, take up valuable space in the operating room, make surgical team coordination challenging, and their proprietary nature makes it difficult to take advantage of recent technological advances, especially in virtual and augmented reality.
+One potential area for further improvement is the integration of modern sensory gloves into robotic platforms, allowing surgeons to control robotic arms intuitively with their hand movements.
+We propose one such system that combines an HTC Vive tracker, a Manus Meta Prime 3 XR sensory glove, and SCOPEYE wireless smart glasses.
 The system controls one arm of a da Vinci surgical robot.
@@
 Hand gestures are used to implement clutching and similar functions.
-In particular, we introduce clutching of the instrument orientation, a functionality not available in the da Vinci system.
+In particular, we introduce clutching of the instrument orientation, a functionality unavailable in the da Vinci system.
 The vibrotactile elements of the glove are used to provide feedback to the user when gesture commands are invoked.
-A preliminary evaluation of the system shows that it has excellent tracking accuracy and allows surgeons to efficiently perform common surgical training tasks with minimal practice with the new interface; this suggests that the interface is highly intuitive.
-The proposed system is inexpensive, allows rapid prototyping, and opens opportunities for further innovations in the design of surgical robot interfaces.
+A qualitative and quantitative evaluation has been conducted that compares the current device with the dVRK console.
+The system is shown to have excellent tracking accuracy, and the new interface allows surgeons to perform common surgical training tasks with minimal practice efficiently.
 I INTRODUCTION
@@
 For the surgeon, it offers increased precision and dexterity through efficient interaction interfaces.
-However, despite its high acceptance and success, it faces some challenges in modern operating rooms.
+However, despite its high acceptance and success, it faces challenges in modern operating rooms (ORs).
 The surgeon consoles are bulky and take up valuable space [1].
@@
 Another challenge with current consoles is that the surgeon is separated from the rest of the team, which can lead to challenges in communication [5] and can impact the efficiency of the operational workflow in tasks such as robotic docking [6].
-Finally, the proprietary nature of existing consoles makes it difficult to innovate in this space despite rapid advances in fields such as augmented and virtual reality that could benefit the surgeon.
+Finally, the proprietary nature of existing consoles makes it challenging to innovate in this space despite rapid advances in fields such as augmented and virtual reality that could benefit the surgeon.
 Several studies have attempted to address some of these shortcomings.
-[7] proposes a novel method for controlling the endoscope through a head-mounted master interface (HMI) integrated with a surgical robot system.
-[8] controls the surgical robot arm and the Augmented Reality (AR) interface through various hand gestures.
-In [5], the authors provide a nice review of existing approaches and opportunities.
-One potential area for further improvement and technological innovation in robotic surgery is the integration of modern sensory gloves into robotic platforms, which allows surgeons to intuitively control robotic arms with their hand movements without the need for additional manipulators that are customary for existing consoles.
+In [7], a novel method is proposed to control the endoscope through a head-mounted interface (HMI) integrated with a surgical robot system.
+And [8, 9] control the surgical robot arm and the Augmented Reality (AR) interface through various hand gestures.
+One potential area for further improvement and technological innovation in robotic surgery is the integration of modern sensory gloves into robotic platforms, which allow surgeons to intuitively control robotic arms with their hand movements without the need for additional manipulators that are customary for existing consoles.
 Furthermore, most end-effectors of surgical robot instruments, such as the Fenestrated Bipolar Forceps (FBF) and the Large Needle Driver (LND), are limited to the use of opposing blades due to the design of the console manipulators (as the last joint is controlled only by pinching two fingers); glove-based control could offer the surgeon additional dexterity while using these instruments.
@@
 These studies explore various sensor modalities and develop communication techniques and control algorithms to achieve sophisticated and effective glove-based control systems for robotic arms.
-In our work, we develop a system in which an HTC Vive Tracker [12] is attached to a Manus Meta Prime 3 Haptic XR sensory glove [13], equipped with multiple IMUs and bending sensors for finger tracking, to control one arm of the da Vinci surgical robot paired with the da Vinci Research Kit (dVRK) [14]; the visual feedback to the surgeon is provided through God Vision wireless smart glasses [15].
+In our work, we develop a system in which an HTC Vive Tracker [13] is attached to a Manus Meta Prime 3 Haptic XR sensory glove [14], equipped with multiple IMUs and bending sensors for finger tracking to control one arm of the da Vinci surgical robot paired with the da Vinci Research Kit (dVRK) [15].
+Visual feedback to the surgeon is provided through SCOPEYE wireless smart glasses [16].
 In addition, this setup manipulates the robotic arm’s end effector (jaw of the instrument) through finger movements.
-Moreover, we implemented hand gesture recognition to substitute pedal inputs of the current da Vinci console (e.g.
-clutch).
-This also allows us to clutch the instrument orientation, a functionality not provided by the da Vinci system.
-The glove is equipped with vibrotactile elements to provide feedback to the user about whether the gesture is activated.
-We evaluate the proposed system in terms of tracking accuracy and the ability of the user to perform surgical training tasks similar to those used in training with the da Vinci devices.
-The evaluation shows that the system is responsive and that surgeons can effectively complete training exercises with minimal practice with the new interface, suggesting that the interface is highly intuitive.
-The proposed system is inexpensive as it uses off-the-shelf components and it dramatically reduces space requirements.
-It demonstrates how modalities such as vibrotactile feedback can efficiently provide additional information to the surgeon without cluttering the visual field.
+Moreover, we implemented hand gesture recognition to substitute pedal inputs of the current da Vinci console (e.g., clutch).
+One of the gestures is employed to clutch the instrument orientation, a functionality not provided by the da Vinci system.
+The glove has vibrotactile sensors to inform the user that a gesture has been activated.
+We evaluated the proposed system by having five surgeons and two medical doctors (no completed residency in surgery) perform a modified version of a robot-assisted surgical exercise using our system and the dVRK console.
+The system has been evaluated in terms of tracking accuracy and movement efficiency.
+After completing the exercise, the participants completed the System Usability Scale [17] and the NASA Task Load Index [18] questionnaires to measure the usability and workload of our proposed architecture.
+These evaluations show that the system is responsive and that surgeons can effectively complete training exercises with minimal practice with the new interface, suggesting that the interface is highly intuitive.
+The proposed system is inexpensive, as it uses off-the-shelf components, and it dramatically reduces space requirements.
+Our work also demonstrates how modalities such as vibrotactile feedback can efficiently provide additional information to the surgeon without cluttering the visual field.
 Finally, the proposed architecture allows for rapid prototyping and opens opportunities for further innovations in the design of surgical robot interfaces.
@@
 Figure 1: The setup showing the configuration of the da Vinci arm (PSM) and the Manus glove with the HTC Vive Tracker attached.
-The user wears the God Vision smart glasses for the stereo endoscopic view.
+The user wears the SCOPEYE smart glasses for the stereo endoscopic view.
 Shown are the $x$ (red), $y$ (green), and $z$ (blue) axes of the relevant coordinate frames.
@@
 The system integrates several components.
-The surgical robot platform consists of the first-generation da Vinci robot paired with the dVRK.
+The surgical robot platform comprises the first-generation da Vinci robot paired with the dVRK.
 One of the Patient Side Manipulators (PSMs) is equipped with the Fenestrated Bipolar Forceps (FBF) instrument and is subject to glove control.
@@
 1.
-The HTC Vive Tracker attached to the Manus glove provides the pose of the hand, while the fingers are tracked by the Manus Meta Prime 3 XR.
-In addition, classifier models were trained for hand gesture recognition; hand gestures were used to supersede da Vinci’s pedal inputs.
-The visual feedback to the user is provided through God Vision smart glasses.
+The HTC Vive Tracker attached to the Manus Meta Prime 3 XR glove provides the pose of the hand, while the fingers are tracked by the glove.
+Instead of using image-based hand-tracking devices as in [9], we decided to use sensor gloves because the hand-tracking performance decreases in the presence of visual occlusion or changes in room lighting.
+These occur frequently in the OR, such as when other surgical team members use secondary screens to look at details in the scene.
+Moreover, sensory gloves offer vibrotactile feedback.
+In addition, classifier models were trained for hand gesture recognition; hand gestures were implemented to substitute the da Vinci console’s pedal inputs.
+SCOPEYE smart glasses that provide the same visual feedback as the High-Resolution Stereo Video (HRSV) of the da Vinci console and have a maximum latency of 7ms [16] were used to further simplify the setup.
 The system runs the Robot Operating System (ROS) [19] that facilitates seamless communication and coordination between the Manus glove, the HTC Vive Tracker, and dVRK, ensuring real-time data streaming and responsive interaction between the user and the control interface.
-The architecture allows for easy prototyping.
 II-B da Vinci Arm Kinematics
 For the full da Vinci system, the kinematic chain from the Patient Side Manipulator (PSM) tip to the Endoscope Camera Manipulator (ECM) tip involves Setup Joints (SUJs).
-The poses provided by the dVRK exhibit errors within a range of ±$\pm 5cm$ for translation and fall between $5\sim 10$ degrees for orientation These inaccuracies arise due to challenges in calibrating the SUJs precisely, as reported in [14].
-Consequently, we implemented a custom calibration method for the dVRK using fiducial markers [17] that eliminates the influence of SUJs and allows us to accurately determine the transformation between the PSM base (frame $R$) and the instrument tip (frame $T$) (Fig.
-1).
-II-C Glove Kinematics
-To monitor the glove’s position and orientation within the environment, we attached an HTC Vive Tracker to it.
-The glove’s position and orientation data are updated at 120 Hz using the data from the IMUs and HTC Vive Base Stations [18, 19].
-The coordinate frames for the HTC Vive Tracker and Base Station are depicted in Fig.
+The poses provided by the dVRK exhibit errors within a range of ±$\pm 5cm$ for translation and fall between $5\sim 10$ degrees for orientation.
+These inaccuracies arise due to challenges in calibrating the SUJs precisely, as reported in [15].
+Consequently, we implemented a custom calibration method for the dVRK using fiducial markers [20] that eliminates the influence of SUJs and allows us to accurately determine the transformation between the PSM base (frame $R$) and the instrument tip (frame $T$), see Fig.
+1.
+II-C Scaling Workspace
+To ensure safety for both the glove and the instrument tip, we define their operational workspaces as cubes, with a larger side $L_{h}$ (30cm) for the glove and a side $L_{t}$ (7.5cm) for the instrument tip, centered at their initial positions.
+In both cases, movements beyond the cube volume are projected to the cube surface.
+The glove movements tend to be significantly larger than the desired movements of the PSM tip ($L_{h}>L_{t}$). The motion of the glove is thus scaled to match the range needed for PSM operations:
+$\Delta p_{t}=\eta\,\Delta p_{h}$ (1)
+where $\Delta p_{h}$ represents the translational movement of the glove from its initial position, $\Delta p_{t}$ denotes the expected translational movement of the PSM from its initial position, and $\eta={L_{t}}/{L_{h}}$ is the scaling factor.
+Similarly to the existing surgeon console, the scaling factor $\eta$ can be adjusted.
+II-D Glove Kinematics
+To monitor the glove’s position and orientation within the environment, we attached a HTC Vive Tracker to it.
+The glove position and orientation data are updated at 120 Hz using the data from the tracker’s IMUs and HTC Vive Base Stations [21, 22].
+Tracking of the relative transformation between the fingers and the wrist is accomplished using the five sets of flexible sensors with 2 Degrees of Freedom (DoF) and six sets of 9 DoF IMUs of the Manus glove, operating at 90Hz with ±$\pm 2.5^{\circ}$ accuracy [14].
+The glove software provides a hand skeleton mesh inspired by the Mediapipe framework [23].
+This mesh is dynamically adjusted based on the readings from the glove sensors, allowing for a real-time representation of the hand pose.
+Hand Gesture Representation Hand Gesture Representation
+Ring [Uncaptioned image] Clutch [Uncaptioned image]
+TABLE I: Gestures trained for distinct functionalities.
+Clutch invokes the clutch mechanism and Ring turns the robot control on or off.
+If neither of these two gestures is recognized, the system returns None.
+The HTC Vive Tracker and Base Station coordinate frames are depicted in Fig.
 1.
 At initialization, the tracker frame $H$ and the tracker base frame $S$ coincide and are aligned with the PSM instrument tip frame $T$.
-A displacement of $H$ from $S$ results in the scaled motion of PSM (see below) to create an intuitive control for the user.
-II-D Control Architecture
-To ensure safety for both the glove and the instrument tip, we define their operational workspaces to be cubes, with a larger side $L_{h}$ for the glove and a side $L_{t}$ for the instrument tip, centered at their initial positions.
-In both cases, movements beyond the cube volume are projected to the cube surface.
-The movements of the glove tend to be significantly larger than the desired movements of the PSM tip ($L_{h}>L_{t}$).
-The motion of the glove is thus scaled to match the range needed for PSM operations:
-$\Delta p_{t}=\eta\,\Delta p_{h}$ (1)
-where $\Delta p_{h}$ represents the translational movement of the glove from its initial position, and $\Delta p_{t}$ denotes the expected translational movement of the PSM from its initial position, and $\eta={L_{t}}/{L_{h}}$ is the scaling factor.
-Similarly to the existing surgeon console, the scaling factor $\eta$ can be adjusted.
-II-E Finger Tracking
-The tracking of the relative transformation between the fingers and the wrist is accomplished using the five sets of flexible sensors with 2 Degrees of Freedom (DoF) and six sets of 9 DoF IMUs of the Manus Meta Prime 3 Haptic XR glove, operating at a rate of 90Hz.
-These sensors boast a ±$\pm 2.5^{\circ}$ accuracy [13].
-The glove software provides a hand skeleton mesh, inspired by the Mediapipe framework [20].
-This mesh is dynamically adjusted based on the readings from the glove sensors, allowing for a real-time representation of the hand pose.
-In our architecture, the Euclidean distance between the tips of the index and thumb fingers is used to control the jaw of the end effector.
-To map the distance of the fingers to the distance between the tips of the jaw we use an expression analogous to Eq.
+The desired translation for the PSM end-effector is calculated by scaling the displacement of $H$ from $S$ as described above to create an intuitive control for the user.
+For the goal orientation of the PSM end effector, we determine the relative orientation between the metacarpophalangeal (MP) and distal interphalangeal (DIP) joint of the sensory glove’s index finger and add this rotation to the one provided by the HTC Vive Tracker readings.
+This ensures that the system follows the index finger of the user, making the system more intuitive and reducing user fatigue.
+In our architecture, the Euclidean distance between the tips of the index and thumb fingers was used to control the jaw of the end effector.
+To map the distance of the fingers to the distance between the tips of the jaw, we use an expression analogous to Eq.
 (1).
-This enables precise and intuitive control of the robotic system.
-II-F Hand Gesture Recognition
-Hand gestures serve as a seamless and intuitive means of interaction and open up the possibility of introducing a large variety of inputs that can surpass what the current consoles can provide.
-For example, in [8] it was shown that it is possible to implement 8 different commands.
+The maximum opening of the jaw is achieved when the distance between the two fingertips is $10$cm.
+This value has been selected as it is the $95\%$ percentile of the length of the index finger of women [24], ensuring that all users can precisely and intuitively control the instrument jaw within the full range of motion.
+II-E Hand Gesture Recognition
+Hand gestures serve as a seamless and intuitive means of interaction and open up the possibility of introducing a wide variety of inputs surpassing what the current consoles can provide.
+For example, [8] shows that it is possible to implement 8 different commands.
 In our work, we train 4 different multiclassification models for hand gesture recognition using our custom dataset.
-Our implemented architecture can currently distinguish between five distinct gestures: None, Pinky (thumb and pinky in contact), Ring (ring finger and thumb in contact), Fist, and Thumb up, as illustrated in Table I.
-Our framework allows a new gesture to be added in just a few minutes.
-Currently, only three of the five gestures serve as commands (Fist, Ring, and Pinky).
-Fist activates the clutch mechanism, Ring, if held for 2 seconds, initiates and suspends robot tracking, and Pinky triggers monopolar current delivery [21].
-See Section III for further details.
-II-G Clutch
+Our implemented architecture can currently distinguish between three distinct gestures: None, Ring (ring finger and thumb in contact), and Clutch (middle, ring, and pinky fingers closed) as illustrated in Table I.
+Clutch activates the clutch mechanism described in Section II-F.
+Holding the Ring gesture for 2 seconds initiates/suspends robot tracking.
+This implementation is open to allow new gestures to be added with a relatively small dataset.
+Further details are provided in Section III.
+II-F Clutch
+Input Hand Pose Predicted Gesture Input Hand Pose Predicted Gesture
+[Uncaptioned image] Ring [Uncaptioned image] Clutch
+TABLE II: Examples of point clouds used for gesture prediction.
 The clutch functionality in the da Vinci surgical robot serves as a safety feature and a means of control for the surgeon during robotic-assisted surgeries.
-This mechanism essentially allows the surgeon to recenter the workspace of the surgeon console by temporarily disengaging the movement of the robotic arms while still maintaining control of the Master Tool Manipulators.
-When the surgeon operates the da Vinci system’s master console, which consists of hand and finger controls, they can engage or disengage the clutch as needed.
-This allows them to stop the movement of the robotic arms, enabling them to make adjustments or reposition the controllers as required during the procedure.
-In the dVRK system, the clutch is engaged with a foot pedal or a button accessible by the index finger.
-However, the current rotation of the tool is maintained and cannot be changed (this is achieved by blocking the last two joints of the controller).
-In contrast, our system implements full clutching (both translation and rotation).
-The clutch is engaged by closing the hand (Fist gesture) as described next.
+This mechanism essentially allows the surgeon to recenter the workspace of the surgeon console by temporarily disengaging the movement of the robotic arms while still maintaining control of the Master Tool Manipulators (MTMs).
+In the dVRK system, the clutch is engaged with a foot pedal.
+However, the current rotation of the tool is maintained and cannot be changed by fixing the orientation of the controllers.
+Our system implements full clutching (both translation and rotation).
+The clutch is engaged by performing the Clutch gesture as described next.
 The clutching mechanism includes 3 phases:
-Leading Edge: When the Fist gesture is detected, the clutch is engaged, and the vibrotactile feedback of the glove is activated.
-The current configuration of the PSM $g_{rt}^{0}$ is recorded and the robot is prevented from moving.
-Note that the jaw of the instrument is allowed to move.
-Active Clutch: The vibrotactile feedback continues to be activated. The PSM is held immobile. The glove can be moved to an arbitrary pose, allowing the user to move to a comfortable configuration.
-Trailing Edge: When the gesture becomes other than Fist, the clutch is disengaged and the vibrotactile feedback is deactivated.
-The tracker frame $H$ and the tracker base frame $S$ are aligned, and their axes are aligned with the axes of the PSM tip frame $T$ whose configuration is described by $g_{rt}^{0}$.
+  • Leading Edge: When the Clutch gesture is detected, the clutch is engaged, activating the vibrotactile feedback of the glove.
+The current configuration of the PSM $g_{rt}^{0}$ is recorded, and the robot is prevented from moving.
+  • Active Clutch: The vibrotactile feedback continues, and the PSM is held steady. The user can change the pose of the glove to a comfortable configuration.
+  • Trailing Edge: When the gesture changes, the clutch is disengaged, and the vibrotactile feedback is deactivated.
+The tracker frame $H$ and the tracker base frame $S$ are aligned to the axes of the PSM tip frame $T$, whose configuration is described by $g_{rt}^{0}$.
 III Hand Gesture Recognition
-Hand Gesture Representation Hand Gesture Representation
-Thumbs up [Uncaptioned image] Pinky [Uncaptioned image]
-Ring [Uncaptioned image] Fist [Uncaptioned image]
-TABLE I: Gestures trained for distinct functionalities.
-In this paper, Fist invokes the clutch mechanism and Ring toggles the robot control on or off.
-None gesture corresponds to the case when the subject is not using any of the four gestures above.
-To train the classifiers for hand gesture recognition, we created a custom dataset.
-We used 5 sets of approximately $3-5$ minute recordings from the glove worn by one of the authors while performing one of the 4 gestures reported in Table I, or none of them.
-The recorded data (Table II) consists of the point cloud of each of the 21 landmark relative poses (3 position entries and 4 quaternion entries) regarding the wrist IMU (147 features).
-Table III shows examples of point clouds for different gestures.
-Each recording has been made while changing the position and orientation of the hand as much as possible to ensure the widest variety of data.
-Each recording corresponds to one gesture (or no gesture) so when merging all the data it was easy to attribute a class to each sample.
-As the number of classes is small, a simple one-hot encoding has been used.
-Hand Gesture None Pinky Ring Fist Thumbs up
-Train set $20743$ $12833$ $25011$ $16738$ $19447$
-Test set $4167$ $2493$ $5075$ $3445$ $3775$
-TABLE II: The size of our training and testing sets for hand gesture recognition.
-Input Hand Pose Predicted Gesture Input Hand Pose Predicted Gesture
-[Uncaptioned image] Thumbs-up [Uncaptioned image] Pinky
-[Uncaptioned image] Ring [Uncaptioned image] Fist
-TABLE III: Examples of point clouds used for gesture prediction.
+We created a custom dataset for training hand gesture recognition, recording approximately $3-5$ minutes of the glove worn by different users performing one of the gestures in Table I, including the None.
+The data consists of point clouds with 21 hand joint landmarks, capturing the relative orientations to the base, set when the glove is activated.
+The orientations are represented by quaternions with respect to the glove base, initialized when the glove is turned on, resulting in a total of 84 features.
+The translations were not added because they vary for the same hand posture with each launch.
+Using only rotations thus makes it easy to capture gestures of interest with only a few recordings.
+Table II shows examples of point clouds for different gestures.
+Each recording contains one gesture, and the pose of the hand is changed to be as diverse as possible.
+We added $150,000$ gestures with sets of random normalized quaternions to the None class to prevent the models from recognizing a random gesture as one of the gestures of interest (Ring or Fist).
+The final dataset consists of $15,557$ Fist instances, $17,036$ Ring instances, and $178,217$ None instances.
+A simple one-hot encoding was applied due to the small number of classes.
 III-A Classifiers
 We initially used Random Forests (RF) [25] as a model for gesture recognition.
-Subsequently, we explored the Multilayer Perceptron (MLP) classifier [23].
-The MLP model was configured with two hidden layers containing 40 and 25 neurons, respectively, and a maximum of 100 iterations.
-The use of such a small model with only two hidden layers helps prevent overfitting.
-In addition to these established models, we investigated a novel classifier, the Light Gradient Boosting Machine (L-GBM) [24].
-L-GBM employs a sequential approach in which decision trees are added successively, with each tree correcting errors from the ensemble of preceding trees.
-Unlike traditional gradient-boosting methods, L-GBM directly minimizes the loss function.
-Tailored for efficient training on large datasets, it is particularly suitable for applications with resource constraints, and it can quickly predict data.
-We also trained a Histogram-based Gradient Boosting model that combines the traditional gradient boosting [25] and the histogram-based algorithm [24].
-Unlike traditional gradient boosting methods that operate on individual data points, histogram-based gradient boosting works with binned feature values.
-By constructing histograms of input features, this approach accelerates the computation of information gained during tree construction.
+Subsequently, we explored the Feedforward Neural Network (FNN) [26] classifiers analogous to those introduced in [27].
+However, the networks have been specifically tuned for classification.
+The structure of the FNN consists of an input layer ($n_{0}$) sized according to the input characteristics, followed by two fully connected layers ($n_{1}$, $n_{2}$), where the number of neurons is proportional to the size of the previous layer ($n_{1}=2\cdot n_{0}$, $n_{2}=2/3\cdot n_{1}$).
+The final layer ($n_{f}$) has $3$ output to function as a classifier, predicting one of the three possible classes.
+In addition to these established models, we investigated the Light Gradient Boosting Machine (L-GBM) [28].Unlike traditional gradient-boosting methods like XGBoost [29], where trees grow depth-wise, L-GBM trees grow leaf-wise, making training faster.
+Tailored for efficient training on large datasets, they are particularly suitable for applications with resource constraints and can quickly provide good predictions.
+We also trained a Histogram-Based Gradient Boosting (HGBT) model that combines traditional gradient boosting [30] and a histogram-based algorithm [28].
+The HGBT works with binned feature values rather than individual data points, which is the case with traditional gradient boosting methods.
+This approach accelerates the computation of information gained during tree construction by constructing histograms of input features.
 The binning process reduces the number of unique feature values, resulting in faster training times and decreased memory usage.
-III-B Results
-By training the models above, it is possible to recognize 5 different gestures (tested on a new, unseen dataset).
-The models were implemented in ROS and Table IV shows their performance.
-The confusion matrices for each model are shown in Fig.
-2.
+To optimize the model hyperparameters, we adopted Grid-Search Cross-Validation (GridSearchCV) [31].
+First, we split the dataset into training and testing sets with the $8:2$ ratio.
+The search process used a 5-fold cross-validation strategy, in which each combination was evaluated across different data splits.
+The grid includes five hyperparameters with three different values, resulting in 1215 combinations after cross-validation.
+The models were evaluated on accuracy, and the best-performing combination of hyperparameters was selected.
+III-B Training Results
 Model Accuracy F1 Score Recall
-MLP 0.9852 0.9852 0.9852
-L-GBM 0.9508 0.9505 0.9508
-RF 0.9417 0.9413 0.9417
-HGBT 0.9628 0.9627 0.9628
-TABLE IV: Evaluation metrics for the trained models.
-Although all models exhibited high accuracy, F1 scores, and recall, we selected the MLP model for implementation due to its high prediction speed (Table V) and high accuracy.
-The average frequency at which the MLP model can operate (determined from the ROS logs) is around 96Hz.
-From Fig.
-2, it is noticeable that the Pinky and Ring gestures are associated with more errors.
-This observation can be attributed to the similarity between these two gestures.
-(a) MLP
-(b) L-GBM
-(c) Random Forest
-(d) HGBT
-Figure 2: Confusion matrices for gesture recognition models.
+FNN 0.9857 0.9764 0.9784
+L-GBM 0.9875 0.9873 0.9875
+RF 0.9830 0.9826 0.9830
+HGBT 0.9792 0.9789 0.9792
+TABLE III: Evaluation metrics of the trained models.
+Table III shows the evaluation results on the test set.
+Based on the metrics, all models had analogous performance in estimating the hand gestures.
+Hence, we conducted an additional live test in which the models predicted the gestures in real time for 1000 iterations each.
+As shown in Table IV, L-GBM was the fastest among the models with an average delay of 1.63 ms.
+The speed is mainly due to its efficient way of growing trees and making their structure as compact as possible.
+The hyperparameters that were used for training the current L-GBM model are as follows: learning_rate=0.1, max_depth=15, min_child_samples=10, n_estimators=100, and num_leaves=15.
+When using the classifiers in real-time applications, the classifier’s output is often post-processed using a majority voting technique [32].
+While this introduces fixed time delays depending on voting buffer size, it enhances the accuracy and reduces short-term false transitions.
+For our application, we set the voting buffer size to 20.
+Thus, the predictions of the L-GBM model over the last 20 iterations are stored in a $20\times 3$ matrix, where each row contains the classifier estimations for each gesture.
+The gesture (column) with the highest sum is selected at the end.
 Model Mean Time (ms) Std. Deviation (ms)
-MLP 0.75 0.14
-L-GBM 2.09 3.06
-RF 0.92 0.16
-HGBT 7.81 4.72
-TABLE V: Prediction time for each model (1000 trials).
-In our application, gesture recognition needs to be performed continuously and in real-time.
-To ensure the stability of recognition – for example, even a single mistake in the clutch could lead to problematic input sent to the robot – we use a sliding window of length 7.
-The predictions of the MLP model over the last 7 time steps are stored in a 7x5 matrix, with each row corresponding to a time sample and each column representing a gesture prediction (0 or 1).
-The matrix is initialized with zeros.
-During each iteration, a new row is added with the most probable gesture set to one; the oldest prediction is deleted from the matrix.
-Finally, in each iteration, the system selects the gesture with the highest column sum.
-This induces a negligible delay but greatly increases the stability of the prediction.
-IV Preliminary Evaluation
-To evaluate the system, the members of the team with surgical training were asked to use the system to complete one of the standard da Vinci training exercises [26].
-The original exercise was modified to suit our system’s single-arm setup and to simplify the task, thereby reducing any uncertainties in the preliminary test.
+FNN 2.70 0.95
+L-GBM 1.63 1.59
+RF 3.90 0.62
+HGBT 4.54 1.40
+TABLE IV: Real-time prediction time for each model (1000 trials).
+IV Experimental Results
+To evaluate our system, 5 surgeons and 2 medical doctors with experience in using the da Vinci surgical console for surgical tasks were asked to use the system to complete one of the standard da Vinci training exercises [33] with both the console and the sensory gloves.
+The original exercise was modified to suit our system’s single-arm setup.
 The modified task required participants to control the PSM with their left hand to lift a peg from one pin on the pegboard and transfer it to another pin.
-One trial consists of repeating this task twice with two different pegs on the pegboard.
-IV-A Experiment Setup
+One trial repeats this task twice with two different pegs on the pegboard.
+IV-A Experimental Setup
 The experimental setup is shown in Fig.
@@
 The endoscope was positioned to allow a view of the pegs on the board and the tip of the instrument.
-God Vision glasses were connected to the vision tower, and the user wore the glasses to access the stereo endoscopic view.
-The evaluation was carried out by 5 surgeons and one fellow, all members of the research team.
-All have prior experience with the da Vinci surgical console to perform surgical tasks.
-Participants underwent a brief training session, lasting approximately 2 to 10 minutes, primarily to familiarize themselves with the operation of the clutch mechanism.
-Each participant performed the exercise 5 times.
-To implement our system, we needed to update the dVRK.
-This caused compatibility issues with the existing dVRK console so we were not able to conduct the exercise through the standard da Vinci console.
+The SCOPEYE smart glasses were connected to the vision tower, providing a stereo endoscopic view to the user.
+Participants underwent a brief training session, lasting approximately 2 to 10 minutes, primarily to familiarize themselves with the glove and the operation of the clutch mechanism.
+Each participant performed the exercise 5 times, both with the glove system and with the dVRK console.
 IV-B Control Results
@@
 (b) Orientation
-(c) Jaw Movements
-Figure 3: A sample recording of motion between the glove and the PSM.
-The segments highlighted in purple indicate periods when the clutch mechanism was engaged.
-(a) displays the translation of the glove and PSM while (b) shows the orientation of the glove and PSM respectively.
-(c) compares the distance of the thumb and index finger on the glove with the jaw angle of the PSM.
-Figure 4: Translational and orientational distance between the glove and the PSM: the left figure depicts the Euclidean distance, while the right one details the orientation difference, quantified with rotation matrices (in the hand workspace).
-Figure 5: Both plots detail the x-axis translation of the glove and the PSM tip in Fig.
-2(a).
-The initial plot illustrates a delay in the robot’s execution of the glove’s movements.
-To address this, the delay between the signals was measured based on the peaks, and the aligned signals were shown on the bottom.
-A sample trajectory of the da Vinci arm controlled with our proposed glove, performed by one of the surgeons, is shown in Fig.
-3.
+(c) Jaw Movement
+Figure 2: A sample trajectory of the glove and the PSM: (a) translation; (b) orientation; and (c) the distance of the thumb and index fingers on the glove and the jaw angle of the PSM.
+The gray and purple regions correspond to the periods when Ring and Clutch gestures, respectively, are active.
+A sample trajectory of the da Vinci arm controlled by one of the surgeons with our proposed system is shown in Fig.
+2.
 A comparison of the kinematics of the glove and the instrument tip is shown in Figs.
 2(a) and 2(b), respectively.
-The motion of the PSM exhibits a stair-case movement due to the implementation of position control of the arm in dVRK (green box in Fig.
-5), where the robot temporarily stops once it reaches the goal position.
-The shaded regions show the moments when the clutch (Fist gesture) was activated.
-Our clutch mechanism (Section II-G) prevents the PSM from moving while activated and resumes the movement smoothly when released.
+The motion of the PSM exhibits a staircase movement due to the implementation of position control of the arm in dVRK (green box in Fig.
+3), where the robot stops once it reaches the reference position.
+In Fig.
+2, the shaded region in gray (3 seconds) corresponds to the initialization of the system.
+This occurs after the Ring gesture has been activated for two seconds.
+During initialization, the arm tries to align with the pose of the glove.
+The shaded region in purple corresponds to the period during which the Clutch was activated.
+The Clutch prevents the PSM from moving; the motion is resumed when the Clutch is released.
 The motion of the index and thumb of the glove and the instrument jaw is depicted in Fig.
 2(c).
-Currently, the fully closed configuration of the FBF jaw is represented by a negative value in the dVRK’s system (as the $0^{\circ}$ angle corresponds to a parallel configuration of the jaw where it is not completely closed); the slight difference for extreme values between the glove and the PSM is due to the scaling between the glove and the jaw.
-To compare the input motion of the glove and the output motion of the instrument tip, we first reversed the scaling (the inverse of Eq.
-(1)) of the movement of the instrument tip (Fig.
-5).
-The figure clearly shows a delay between the two trajectories, mainly due to the computation time for the inverse kinematics.
-To accurately estimate the error between the input and output signals and eliminate the effect of the delay, we identified the delay by the peak positions within each signal.
-Subsequently, the PSM signal is shifted according to the calculated delay to align with the glove’s signal (Fig.
-5).
-The difference in translation was quantified using the Euclidean distance, while the discrepancy in rotation was determined using the standard bi-invariant metric on the group of rotations [27].
-The errors derived from Figs.
-2(a) and 2(b) are shown in Fig.
-4.
-The duration and error metrics for each participant’s performance are depicted in Table VI.
-The average delay in all trials and all participants was 0.223s.
-User Average Duration (s)
+Figure 3: The x-axis translation of the glove and the PSM tip: raw data (top) and data after the delay has been eliminated (bottom).
+Controller
 Translational
@@
 Rotational
-Error (rad)
-Average Std. dev. Average Std. dev.
-A 60.345 0.005 0.004 0.024 0.024
-B 45.304 0.004 0.004 0.016 0.018
-C 57.922 0.005 0.004 0.040 0.056
-D 53.298 0.004 0.004 0.021 0.026
-E 52.506 0.003 0.003 0.015 0.015
-F 37.782 0.003 0.003 0.021 0.021
-TABLE VI: the first column displays the average runtime for the five trials.
-The subsequent columns detail the overall mean and standard deviation of the errors, calculated based on the glove’s workspace, performed by each participant (complete version available on [28]).
+Error (rads)
+Delay (s)
+Mean Std. Dev. Mean Std. Dev.
+Glove 0.005 0.010 0.096 0.146 0.205
+Console 0.002 0.001 0.038 0.066 0.001
+TABLE V: The average translational and rotational errors between each controller and the PSM across all participants.
+The last column shows the average delay between the controller and the PSM tip signals.
+User Glove Console
+Avg. Time
+Avg. Travel
+Distance (m)
+Avg. Time
+Avg. Travel
+Distance (m)
+A 42.7 0.503 51.9 0.585
+B 40.3 0.541 35.8 0.479
+C 33.5 0.442 29.2 0.375
+D 39.3 0.606 18.5 0.298
+E 50.2 0.726 30.8 0.365
+F 38. 0.479 27.5 0.302
+G 34.3 0.503 27.1 0.382
+Total 39.8 0.543 31.6 0.398
+TABLE VI: Performance evaluation.
+The first column displays the average runtime, and the subsequent column shows the average traveled Euclidean distance of the instrument tip manipulated by each controller of the five trials from each participant.
+To compare the input motion of the glove and the output motion of the instrument tip, we first reversed the scaling (Eq.
+(1)) of the movement of the instrument tip (Fig.
+3).
+The figure shows a delay between the two trajectories, mainly due to the computation time for the inverse kinematics.
+To accurately estimate the error between the input and output signals, we identified the delay by the peak positions within each signal and eliminated the effect of the delay by shifting the PSM signal to align with the glove’s signal (Fig.
+3).
+The difference in translation was quantified using the Euclidean distance, while the discrepancy in rotation was determined using the standard bi-invariant metric on the group of rotations [34].
+The errors for each controller in all the trials across the participants are listed in Table V.
+Although the magnitude of the errors is larger for our system, the task completion times in the two setups were comparable (Table VI), suggesting that the errors were acceptable.
+These larger errors might also explain the slightly higher average travel distances for our system.
+The main reason for significantly higher std.
+deviation for our system is mainly the consequence of safety mechanisms in the dVRK console: when the difference between the PSM position and the console position is larger than a threshold (as when the PSM is held by a peg, for example), the system shuts down.
+No such mechanism is employed in our system, leading to the presence of large errors.
+The average delay of our system is $0.2$s, significantly larger than for the dVRK console. However, [35] shows that such delays are routinely tolerated by humans.
+We evaluated the average performance of the participants with each controller (Table VI).
+Performance is evaluated by the average duration and travel distance of the instrument tip, indicating how efficiently they managed to perform the task.
+Considering that this is a prototype system, the performance of controlling the robot with the glove compared favorably with that of the console.
 IV-C Survey Results
-In the follow-up survey, each participant was asked five questions: 1) How comfortable was the glove controller?
-2) How responsive was the robot when controlled with the glove controller?
-3) How intuitive was controlling the jaw with your fingers?
-4) How intuitive was it to control the arm with the glove?
-5) Was the vibrotactile feedback for the clutch effective?
-We use a 7-point Likert scale for the response, with 7 being the most positive response [29].
-The average responses for the 6 participants are summarized in Table VII.
-The results of Questions 3 to 5 show that the new control architecture is intuitive.
-This is further evidenced by the small amount of practice required to complete the exercise.
-Questions 1 and 2 produced satisfactory results; all participants responded that the system was ’highly responsive.’ However, the results of Question 1 indicate that the overall ergonomics of the system can be improved, as most users had to lift their elbow to achieve a correct angle with their hand.
-This issue also highlights two important aspects: first, the user’s experience with the device may significantly impact their comfort, as they can readjust themselves by clutching to find a better position.
-Second, an ergonomic seat can greatly reduce the user’s effort.
-Question # Q1 Q2 Q3 Q4 Q5
-Avg. Score 5.50 6.00 6.50 6.50 6.33
-TABLE VII: Average scores in the Likert scale from the users.
+Participants were asked to compare the usability and task load of our proposed system with the conventional console.
+Usability was assessed with the System Usability Scale (SUS) [17], and task load was measured using NASA Task Load Index (TLX) [18] questionnaires.
+The overall results are shown in Table VII.
+The average SUS score for the glove was 67.5 compared to the console’s 81.4 (13.9 points of difference).
+These results align well with the quantitative assessment of the performance in Table VI: users who achieved similar performance between the glove and the console responded with a similar SUS assessment for the two.
+Considering that users have significantly more training on the dVRK console and that our system is just a prototype, we judge the reported SUS score as overwhelmingly positive.
+When considering the workload, the average TLX score of the glove was higher than that of the console (40.6 vs.
+26.1).
+We hypothesize that this is mainly due to the lack of ergonomic support in our system, while the dVRK console is equipped with a proper armrest and headrest.
+In addition, as in the case of the SUS scores, users are more familiar with the da Vinci console setup, significantly reducing the effort needed to adapt to the system.
+Thus, we anticipate that these scores would improve if users spent more time with our system.
 V Conclusion
-This paper presents a novel user interface for the da Vinci surgical robot arm.
-The system integrates the Manus Meta Prime 3 Haptic XR glove, the HTC Vive Tracker, God Vision wireless smart glasses, and controls one arm of the dVRK surgical robot platform.
+This paper presents a novel surgeon interface for the da Vinci surgical robot.
+The system integrates the Manus Meta Prime 3 Haptic XR glove, the HTC Vive Tracker, and SCOPEYE wireless smart glasses and controls one arm of the dVRK surgical robot platform.
 In addition to moving the arm, the surgeon can use fingers to control the end-effector of the surgical instrument.
-A custom robot calibration was implemented to address the inherent inaccuracies of the dVRK system.
-A combination of scaling and gesture recognition was used to achieve precise and intuitive control of the robot’s movements while implementing an upgraded version of the clutch mechanism.
-In particular, we introduced clutching of the instrument orientation, a functionality not available in the da Vinci system.
-The vibrotactile elements of the glove are used to provide feedback to the user when gesture commands are invoked.
+A combination of scaling and gesture recognition was used to achieve precise and intuitive control of the robot’s movements while implementing an enhanced version of the clutch mechanism.
+In particular, we introduce clutching of the instrument orientation, a functionality unavailable in the da Vinci system.
+The vibrotactile feedback of the glove is used to alert the user when gesture commands are invoked.
 The proposed interface offers several potential advantages.
-Surgeons can directly control robotic arms by moving their hands, without the need for additional physical devices.
+Surgeons can directly control robotic arms by moving their hands without the need for additional physical devices.
 This facilitates a more immersive surgeon-robot interaction.
-However, it should be noted that the hardware used in our setup is not the latest available on the market; we used the first version of HTC Vive Trackers and Manus Meta Prime 3 XR, both of which have lower precision compared to their upgraded versions (HTC Vive Trackers v2 and Quantum XR Metaglove).
-The system was tested by members of the team with surgical training, including 5 surgeons and 1 fellow.
+However, it should be noted that the hardware used in our setup is not the latest available on the market; we used the first version of HTC Vive Trackers and Manus Meta Prime 3 XR, both of which have lower precision compared to the next generation devices (HTC Vive Trackers v2 and Quantum XR Metaglove).
+Controller
+Glove Console
+User SUS TLX SUS TLX
+A 92.5 27 92.5 35
+B 70 45 85 27
+C 67.5 45 60 65.8
+D 80 10.8 80 5
+E 22.5 60.8 92.5 11.6
+F 82.5 48 92.5 12.5
+G 57.5 47 67.5 25
+Total Average 67.5 40.6 81.4 26.1
+TABLE VII: System Usability Scale (SUS) and NASA Task Load Index (TLX) questionnaire results, answered by surgeons (green letters) and medical doctors (red letters).
+In the SUS, a score closer to 100 indicates better usability, and in the TLX, a score closer to 100 indicates a higher workload.
+The system was tested by users who were experienced with the da Vinci system, including five surgeons and two medical doctors.
 They were asked to perform a surgical training task similar to those used in training with the da Vinci devices.
 The evaluation shows that the system is responsive and precise.
-Surgeons were able to effectively complete training exercises with minimal practice with the new interface, suggesting that the interface is highly intuitive.
-After the exercise, each participant filled out a questionnaire to provide feedback on their user experience.
-The responses show a high level of user satisfaction.
-Existing surgical consoles are bulky and take up valuable space.
+Surgeons could effectively complete training exercises with minimal practice with the new interface, suggesting that the interface is highly intuitive.
+After the exercise, each participant completed two questionnaires to assess the usability and workload of the proposed system.
+Users rated their experience favorably; however, a higher workload suggests the need for a more ergonomic setup.
+The existing surgical consoles are bulky and take up valuable space in the operating room.
 They limit the ability to optimize the operating room setup for throughput.
 They also pose challenges to the communication of the surgical team, leading to potential workflow disruptions.
-In contrast, our proposed system is inexpensive as it uses off-the-shelf components, and it dramatically reduces space requirements.
+In contrast, our proposed system is inexpensive, as it uses off-the-shelf components, and it dramatically reduces space requirements.
 It demonstrates how modalities such as vibrotactile feedback can efficiently provide additional information to the surgeon without cluttering the visual field.
 Finally, the proposed architecture allows for rapid prototyping and opens opportunities for further innovations in the design of surgical robot interfaces.
-Future work includes implementing velocity control to enhance user experience, upgrading hardware to improve tracking precision, enabling full control for the right hand, and expanding the repertoire of user gestures to increase the range of available actions.
+Future work includes implementing velocity control to enhance the user experience, enabling full control of both surgical arms,upgrading hardware to improve tracking precision, and expanding the repertoire of user gestures to increase the range of available actions.

### v3

@@
 One potential area for further improvement and technological innovation in robotic surgery is the integration of modern sensory gloves into robotic platforms, which allow surgeons to intuitively control robotic arms with their hand movements without the need for additional manipulators that are customary for existing consoles.
-Furthermore, most end-effectors of surgical robot instruments, such as the Fenestrated Bipolar Forceps (FBF) and the Large Needle Driver (LND), are limited to the use of opposing blades due to the design of the console manipulators (as the last joint is controlled only by pinching two fingers); glove-based control could offer the surgeon additional dexterity while using these instruments.
+Furthermore, most end-effectors of surgical robot instruments, such as the Fenestrated Bipolar Forceps (FBF) and the Large Needle Driver (LND), are limited to the use of opposing blades  [10] due to the design of the console manipulators (as the last joint is controlled only by pinching two fingers); glove-based control could offer the surgeon additional dexterity while using these instruments.
 Glove-based control is intuitive and is popular in robotics applications outside the surgical field.
@@
 The HTC Vive Tracker attached to the Manus Meta Prime 3 XR glove provides the pose of the hand, while the fingers are tracked by the glove.
-Instead of using image-based hand-tracking devices as in [9], we decided to use sensor gloves because the hand-tracking performance decreases in the presence of visual occlusion or changes in room lighting.
+Instead of using image-based hand-tracking devices as in [9].
+We decided to use sensory gloves because video-based hand-tracking (commonly used in AR, VR, and Leap Motion applications) performs poorly in the presence of visual occlusion, background noise, or variations in room lighting [20].
 These occur frequently in the OR, such as when other surgical team members use secondary screens to look at details in the scene.
@@
 These inaccuracies arise due to challenges in calibrating the SUJs precisely, as reported in [16].
-Consequently, we implemented a custom calibration method for the dVRK using fiducial markers [20] that eliminates the influence of SUJs and allows us to accurately determine the transformation between the PSM base (frame $R$) and the instrument tip (frame $T$), see Fig.
+Consequently, we implemented a custom calibration method for the dVRK using fiducial markers [22] that eliminates the influence of SUJs and allows us to accurately determine the transformation $g_{rt}$ between the PSM base (frame $R$) and the instrument tip (frame $T$), see Fig.
 1.
@@
 [Uncaptioned image] Ring [Uncaptioned image] Clutch
-TABLE II: Examples of point clouds used for gesture prediction.
+TABLE II: Examples of point clouds used as input
 The clutch functionality in the da Vinci surgical robot serves as a safety feature and a means of control for the surgeon during robotic-assisted surgeries.
@@
 III-B Training Results
-Model Accuracy F1 Score Recall
-FNN 0.9857 0.9764 0.9784
-L-GBM 0.9875 0.9873 0.9875
-RF 0.9830 0.9826 0.9830
-HGBT 0.9792 0.9789 0.9792
-TABLE III: Evaluation metrics of the trained models.
+Model Accuracy F1 Score Recall Mean (ms) Std. Dev. (ms)
+FNN 0.9857 0.9764 0.9784 2.70 0.95
+L-GBM 0.9875 0.9873 0.9875 1.63 1.59
+RF 0.9830 0.9826 0.9830 3.90 0.62
+HGBT 0.9792 0.9789 0.9792 4.54 1.40
+TABLE III: Evaluation metrics of the trained models. The last two columns represent the real-time prediction time for each model (estimated over 1000 trials).
 Table III shows the evaluation results on the test set.
@@
 Hence, we conducted an additional live test in which the models predicted the gestures in real time for 1000 iterations each.
-As shown in Table IV, L-GBM was the fastest among the models with an average delay of 1.63 ms.
+As shown in Table III, L-GBM was the fastest among the models with an average delay of 1.63 ms.
 The speed is mainly due to its efficient way of growing trees and making their structure as compact as possible.
@@
 The gesture (column) with the highest sum is selected at the end.
-Model Mean Time (ms) Std. Deviation (ms)
-FNN 2.70 0.95
-L-GBM 1.63 1.59
-RF 3.90 0.62
-HGBT 4.54 1.40
-TABLE IV: Real-time prediction time for each model (1000 trials).
 IV Experimental Results
@@
 Console 0.002 0.001 0.038 0.066 0.001
-TABLE V: The average translational and rotational errors between each controller and the PSM across all participants.
+TABLE IV: The average translational and rotational errors between each controller and the PSM across all participants.
 The last column shows the average delay between the controller and the PSM tip signals.
@@
 Total 39.8 0.543 31.6 0.398
-TABLE VI: Performance evaluation.
+TABLE V: Performance evaluation.
 The first column displays the average runtime, and the subsequent column shows the average traveled Euclidean distance of the instrument tip manipulated by each controller of the five trials from each participant.
@@
 The difference in translation was quantified using the Euclidean distance, while the discrepancy in rotation was determined using the standard bi-invariant metric on the group of rotations [36].
-The errors for each controller in all the trials across the participants are listed in Table V.
-Although the magnitude of the errors is larger for our system, the task completion times in the two setups were comparable (Table VI), suggesting that the errors were acceptable.
+The errors for each controller in all the trials across the participants are listed in Table IV.
+Although the magnitude of the errors is larger for our system, the task completion times in the two setups were comparable (Table V), suggesting that the errors were acceptable.
 These larger errors might also explain the slightly higher average travel distances for our system.
@@
 The average delay of our system is $0.2$s, significantly larger than for the dVRK console. However, [37] shows that such delays are routinely tolerated by humans.
-We evaluated the average performance of the participants with each controller (Table VI).
+We evaluated the average performance of the participants with each controller (Table V).
 Performance is evaluated by the average duration and travel distance of the instrument tip, indicating how efficiently they managed to perform the task.
@@
 Usability was assessed with the System Usability Scale (SUS) [18], and task load was measured using NASA Task Load Index (TLX) [19] questionnaires.
-The overall results are shown in Table VII.
+The overall results are shown in Table VI.
 The average SUS score for the glove was 67.5 compared to the console’s 81.4 (13.9 points of difference).
-These results align well with the quantitative assessment of the performance in Table VI: users who achieved similar performance between the glove and the console responded with a similar SUS assessment for the two.
+These results align well with the quantitative assessment of the performance in Table V: users who achieved similar performance between the glove and the console responded with a similar SUS assessment for the two.
 Considering that users have significantly more training on the dVRK console and that our system is just a prototype, we judge the reported SUS score as overwhelmingly positive.
@@
 Total Average 67.5 40.6 81.4 26.1
-TABLE VII: System Usability Scale (SUS) and NASA Task Load Index (TLX) questionnaire results, answered by surgeons (green letters) and medical doctors (red letters).
+TABLE VI: System Usability Scale (SUS) and NASA Task Load Index (TLX) questionnaire results, answered by surgeons (green letters) and medical doctors (red letters).
 In the SUS, a score closer to 100 indicates better usability, and in the TLX, a score closer to 100 indicates a higher workload.
@@
 They also pose challenges to the communication of the surgical team, leading to potential workflow disruptions.
-In contrast, our proposed system is inexpensive, as it uses off-the-shelf components, and it dramatically reduces space requirements.
+In contrast, our proposed system is inexpensive (estimated total cost of  \$5000), as it uses off-the-shelf components.
+Finally it dramatically reduces space requirements and is easily portable.
+The dimensions of the da Vinci S/Si console are 166 cm x 163 cm x 163 cm and it weighs 273 kg, compared to our setup including 2 HTCVive tracker (7cm × 8cm × 5cm, 100g), 2 basestations (9cm x 6cm x 8cm 0.45 kg ), 2 SCOPEYE wireless input transmitter (in total 39cm x 30cm x 16cm) two sensory gloves (worn by the surgeon, 138g) and a computer.
 It demonstrates how modalities such as vibrotactile feedback can efficiently provide additional information to the surgeon without cluttering the visual field.
@@
 Future work includes implementing velocity control to enhance the user experience, enabling full control of both surgical arms, upgrading hardware to improve tracking precision, and expanding the repertoire of user gestures to increase the range of available actions.
+We will also implement enhanced ergonomic features, such as adjustable armrests and customizable control environments.